# Superpixel Graph Contrastive Clustering with Semantic-Invariant Augmentations for Hyperspectral Images

### v1

@@ V Conclusion
In this article, we proposed a superpixel graph contrastive clustering (SPGCC) algorithm to learn clustering-friendly representations of HSI. In SPGCC, high-order spatial and spectral information of HSI is extracted and preserved by pixel-level pre-training with 3-D and 2-D hybrid CNNs. To conduct superpixel-level contrastive clustering, we design two semantic-invariant superpixel data augmentations, i.e., pixel sampling augmentation and model weight augmentation, where internal pixels and the output of dual-branch GCN serve as augmented views of superpixels. In addition, we recompute the high-confidence center of each class to mitigate the influence of outliers. The optimization target is composed of sample-level alignment and clustering-center-level contrast to overcome the weakness of common contrastive learning that pushes representations in the same class away. Experimental results demonstrated that SPGCC outperforms compared methods on different HSI datasets significantly. In the future, we hope to mine the more accurate correlation between superpixels to further guide superpixel-level graph learning for HSI.

### v2

@@ V Conclusion
In this article, we proposed a superpixel graph contrastive clustering (SPGCC) algorithm to learn clustering-friendly representations of HSI. In SPGCC, high-order spatial and spectral information of HSI is extracted and preserved by pixel-level pre-training with 3-D and 2-D hybrid CNNs. To improve superpixel-level contrastive clustering, we design two semantic-invariant superpixel data augmentations, i.e., pixel sampling augmentation and model weight augmentation, where internal pixels and the output of dual-branch GCN serve as augmented views of superpixels. In addition, we recompute the high-confidence center of each class to mitigate the influence of outliers. Compared with general graph data augmentations, the proposed augmentations are more reliable. Clustering is guided by sample-level alignment and clustering-center-level contrast, which can overcome the weakness of common contrastive learning that pushes samples in the same class away and help learn the representations with intra-class similarity and inter-class dissimilarity. Experimental results demonstrated that SPGCC outperforms compared methods on different HSI datasets significantly. In the future, we hope to mine the more accurate correlation between superpixels to further guide superpixel-level graph learning for HSI.